pretend your group must make 2 M NaCl in a total volume of 100 mL water using the glassware indicated in the procedure. Create a bulleted list of how to make this solution that your partner could follow.

.
Answer #1

First, let us do some calculations to find the mass of NaCl to be taken:

Molarity = moles of solute/volume of solution in L

Moles of solute = 2*0.1 = 0.2 mol

Molar mass of NaCl = 58.44 g/mol

Mass of NaCl to be used = 0.2*58.44 = 11.688 g

Procedure :

  • Take 11.688 g (11.7 g approximately) of NaCl. The mass is calculated by subtracting the mass of empty beaker from the mass of beaker+ NaCl.
  • Measure 100 mL of water using measuring cylinder.
  • Mix the taken NaCl with 100 mL water.

This gives 2 M NaCl in 100 mL water.